OpenStudy (anonymous):

a person flying a kite holds the string 4 feet above ground level. the string of the kite is taut and makes an angle of 60 degrees with the horizontal. Approximate the height of the kite above level ground if 500 feet of string is payed out

OpenStudy (wolf1728):

sine (60°) = hgt / hypotenuse sine (60°) = hgt / 500 feet 0.8660254038 * 500 = height height = 433 feet Adding the 4 feet above the ground where the kite is held height = 437 feet

OpenStudy (anonymous):

\[500 \sin (60 {}^{\circ})+4=4+250 \sqrt{3}=437.013 \]